# Neural Collapse-Inspired Multi-Label Federated Learning under Label-Distribution Skew

### v1

@@ Introduction
In a multi-label, label-skewed FL setting, a total of $K$ clients collaboratively train a shared global model. Each client $k$ has a local dataset $D_{k}=\{(x^{k}_{i},y^{k}_{i})\}_{i=1}^{N_{k}}$, containing $N_{k}$ samples. Each sample $(x^{k}_{i},y^{k}_{i})$ consists of an input image $x^{k}_{i}\in\mathbb{R}^{W\times H\times 3}$ and a corresponding multi-hot label vector $y^{k}_{i}\in\mathbb{R}^{C}$, indicating the presence of one or more of the $C$ possible classes in the image. Due to the label-skewed nature of the scenarios, each client may only have access to a subset of the total class set, denoted as $C_{k}\subseteq C$. Moreover, even for the present classes, the data distribution may differ significantly across clients. The objective of the federated system is to learn a global model that can accurately recognise all $C$ classes, without requiring clients to share their local data.
Neural Collapse
@@
Simplex Equiangular Tight Frame (ETF). A simplex ETF matrix $\mathbf{M}=\{\mathbf{m}_{c}\in\mathbb{R}^{d}\mid c=1,2,\dots,C\}$ composed of $C$ column vectors, each corresponding to a class prototype in $\mathbb{R}^{d}$. The matrix is defined as:
$\mathbf{M}=\sqrt{\frac{C}{C-1}}\mathbf{U}(\mathbf{I}_{C}-\frac{1}{C-1}\mathbf{J}_{C})$ (1)
$\mathbf{m}_{a}^{\top}\mathbf{m}_{b}=\begin{cases}-\frac{1}{C-1},&\text{if }a\neq b\\ 1,&\text{otherwise}\end{cases}$ (2)
where $\mathbf{U}\in\mathbb{R}^{d\times C}$ denotes a rotation orthogonal matrix $\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}_{C}$. $\mathbf{I}_{C}\in\mathbb{R}^{C\times C}$ is the identity matrix. $\mathbf{J}_{C}\in\mathbb{R}^{C\times C}$ is an all-ones matrix. All vectors in the Simplex ETF have equal $\ell_{2}$ norm and identical pair-wise angles.
Variability Collapse ($\mathcal{NC}_{1}$). The intra-class variability of final-layer features collapses to zero, meaning that all feature vectors corresponding to the same class converge to their corresponding class mean.
Convergence to Simplex ETF ($\mathcal{NC}_{2}$). The class means collapse to the vertices of a simplex ETF, forming a maximally symmetric and equidistant configuration.
Self-Duality ($\mathcal{NC}_{3}$). Upon appropriate rescaling, the weights of the final classification layer also collapse to the class means. This results in a self-dual structure, where both the classifier weights and class means share the same simplex ETF structure.
@@
The discovery and analysis of NC in prior work reveal a promising structure for optimal feature clustering in the latent feature space under ideal training conditions. This insight provides valuable guidance on how to shape the latent feature space when data distributions are imperfect. In this work, we aim to leverage the desirable properties of NC within the challenging multi-label FL setting. An overview of the proposed framework is illustrated in Figure 2.
Label-Aware Disentanglement Module
Multi-label classification involves processing images that may contain multiple objects from different categories. This poses challenges for conventional image classification pipelines, which typically employ global average pooling to produce image-level feature representations. However, such pooled representations are often suboptimal for multi-label tasks because they compress multiple semantic concepts into a single vector, making it difficult to learn distinct, class-specific features in the latent feature space. Additionally, the varying number of labels per image further complicates representation learning and limits a model’s ability to structure and control the learned features effectively.
To address these challenges, we draw inspiration from recent advances in multi-label learning (Liu et al.,, 2021; Ridnik et al.,, 2023; Xu et al.,, 2023), which focus on extracting label-specific features. Building on this idea, we incorporate an attention mechanism to extract class-specific information from the image-level representation, thereby enhancing the discriminative power of each label. In our proposed pipeline, each input image is first processed by a standard feature extractor such as a CNN or transformer to produce a global feature representation. This is then passed to the Label-Aware Disentanglement Module (LADM), which utilises a 4-head cross-attention mechanism to derive label-specific features. In the FL setting, each client trains exclusively on its local data, which often deviates significantly from the global distribution. To address this issue, we adopt a predefined ETF matrix, used both as the classifier and as the source of class-wise query embeddings in LADM. By sharing this fixed query matrix across all clients, LADM ensures a consistent strategy for extracting disentangled, class-specific features during local training. This consistency reduces conflicts during model aggregation and promotes better feature alignment in the global model. The following equations detail the overall process:
$\mathbf{H}$ $=\text{LADM}(\mathbf{M},\mathbf{f})$ (3)
$=\text{Concat}_{0}\left(\text{MultiHeadAttn}(\mathbf{m}_{c},\mathbf{f})\right)_{c=1}^{C},$
where $\mathbf{f}\in\mathbb{R}^{d}$ denotes the image-level feature, serving as both the key and value in the attention mechanism. The output $\mathbf{H}\in\mathbb{R}^{C\times d}$ represents the class-specific features generated by LADM. $\mathbf{M}\in\mathbb{R}^{d\times C}$ is a predefined ETF matrix that provides class-wise query embeddings, where $\mathbf{m}_{c}$ is the query for class $c$. $\text{Concat}_{0}$ denotes concatenation along dimension 0.
@@
After disentangling image-level features into class-specific representations, each feature captures a distinct semantic concept. This enables the effective adaptation of established representation learning techniques from the multi-class setting to the multi-label context. Rather than allowing each local model to independently learn its own classification boundaries, which can be hindered by label imbalance or missing classes, we fix the classifier using a predefined simplex ETF matrix. This design enforces maximal angular separation between class embeddings, enhancing the model’s ability to distinguish between classes regardless of the local label distribution. By reducing the influence of local label co-occurrence, this approach also helps mitigate divergence between local and global data distributions. Once class-specific features are extracted via LADM, each feature is projected onto its corresponding ETF classifier weight to compute a class-specific logit. These logits are then passed through a sigmoid activation to obtain binary predictions. The model is trained using binary cross-entropy (BCE) loss. The process is formalised as follows:
$\hat{y}_{i}^{c}=\text{Sigmoid}(\mathbf{h}^{c}_{i}\cdot\mathbf{m}_{c})$ (4)
$\mathcal{L}_{\text{BCE}}=-\frac{1}{C}\frac{1}{N}\sum_{c=1}^{C}\sum_{i=1}^{N}\left[y_{i}^{c}\log(\hat{y}_{i}^{c})+(1-y_{i}^{c})\log(1-\hat{y}_{i}^{c})\right]$ (5)
where $\mathbf{h}_{i}^{c}\in\mathbb{R}^{d}$ denotes the class-specific feature for class $c$ extracted by LADM for the ii‑th sample, and $\mathbf{m}_{c}\in\mathbb{R}^{d}$ represents the fixed ETF classifier weight corresponding to class $c$. $C$ is the total number of classes, and $N$ denotes the number of training samples available on the client.
@@
After extracting class-wise features using LADM, each sample is associated with multiple features, corresponding to the total number of classes. Based on the ground-truth labels, a subset of these features is considered as positive, while the remaining ones are treated as negative. To promote more effective clustering in the latent feature space, we introduce two additional regularization terms. These terms are designed to suppress potential noise introduced by negative features and to encourage compact clustering of positive features belonging to the same class.
Negative Feature Rejection Loss: The $\mathcal{L}_{\text{BCE}}$ loss encourages each class-wise feature to align with its corresponding ETF prototype if it corresponds to a ground-truth (positive) label, and to be pushed away from its prototype if it corresponds to a negative label. To further reduce prediction ambiguity, it is important to prevent negative features from aligning with any other class prototypes. To address this, we compute the similarity between negative features and all class-wise prototypes, applying a penalty to high similarity scores. This regularization helps suppress the influence of negative features on the final prediction and mitigates potential noise. The process is formalised as follows:
$\hat{y}_{ij}^{-}=\text{Sigmoid}(\mathbf{h}^{-}_{i}\cdot\mathbf{m}_{j})$ (6)
$\mathcal{L}_{\text{Neg}}=-\frac{1}{C}\frac{1}{N}\sum_{j=1}^{C}\sum_{i=1}^{N}\mathbb{I}(\hat{y}_{ij}^{-}>\tau)\cdot\log(1-\hat{y}_{ij}^{-})$ (7)
where $h_{i}^{-}\in\mathbb{R}^{d}$ denotes the negative feature of sample $i$. The threshold $\tau$ is a predefined value used to filter out negative features with low similarity scores, ensuring that only confident negative predictions contribute to the loss. In our experiments, we empirically set $\tau=0.3$.
@@
Positive Feature Contrastive Loss: To encourage more compact and discriminative class-wise clustering in the latent feature space, we introduce a contrastive loss. This loss enforces positive features to remain close to their corresponding class prototypes while being pushed away from all other class prototypes. The formulation is:
$\mathcal{L}_{\text{Pos}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{e^{\mathbf{h}^{c}_{i}\cdot\mathbf{m}_{c}}}{\sum_{j=1}^{C}e^{\mathbf{h}^{c}_{i}\cdot\mathbf{m}_{j}}}$ (8)
where $\mathbf{h}^{c}_{i}\in\mathbb{R}^{d}$ denotes the positive feature of sample $i$ for class $c$. $\mathbf{m}_{j}$ represents the predefined prototype for class $j$.
With the two regularization terms, one designed to suppress noise from negative features and the other to encourage compact clustering of positive features, the total loss can be expressed as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{BCE}}+\lambda_{1}\mathcal{L}_{\text{Neg}}+\lambda_{2}\mathcal{L}_{\text{Pos}}$ (9)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters used to balance the BCE loss and the two regularization terms.
@@
(a) FC classifier
(b) LADM + fixed ETF classifier
(c) LADM + ETF + Dual regularization
@@ Acknowledgments
Pseudo Code for FedNCAlign-ML
We present the pseudocode for our proposed FedNCAlign-ML in Algorithm 1. The input comprises a set of clients, each with its own local dataset, and a centralized server responsible for managing communication. The training process involves three iterative steps: local training on client devices, aggregation at the central server, and broadcasting the updated global model back to the clients. While we adopt the standard FL protocol for aggregation and broadcasting, the core contribution of FedNCAlign-ML lies in its specialized model design and the local training procedure tailored for multi-label label-skewed FL scenarios. During local training, each client first encodes its input data using a standard feature extractor to generate image-level representations. These representations are then passed to the proposed Label Aware Disentanglement Module (LADM), which produces disentangled, class-wise embeddings. By capturing distinct semantic concepts within each feature, these embeddings facilitate more effective and controllable clustering in the latent feature space. To further enhance consistent and discriminative feature alignment across clients, particularly under label-skewed conditions, we introduce a shared, predefined Equiangular Tight Frame (ETF) structure. This structure encourages class-independent clustering and aligns local features with a global class geometry, helping to mitigate overfitting to local data distributions. Additionally, two complementary regularization terms are introduced to enhance feature clustering in the latent feature space: a rejection loss to filter out noisy signals from negative features, and a contrastive loss to promote compact intra-class clustering and clear inter-class separation.
Algorithm 1 Federated Neural Collapse Alignment for Multi-Label Learning (FedNCAlign-ML)
Input: $K$ clients; local datasets $\{\mathcal{D}_{k}\}_{k=1}^{K}$; initial global parameters $w_{0}$; predefined ETF matrix $\mathbf{M}$; learning rate $\eta$; local epochs $E$.
1:Server:
2:Initialize $w\leftarrow w_{0}$
3:for $t=0$ to $T$ do $\triangleright$ $t$-th communication round
4:  for each client $k=1,\dots,K$ in parallel do
5:   $w_{t+1}^{k}\leftarrow\textsc{ClientUpdate}(\mathcal{D}_{k},w_{t})$
6:   Send the client model $w^{k}_{t+1}$ to the server
7:  end for
8:  $w_{t+1}\leftarrow\textsc{ModelAggregation}(\{w_{t+1}^{k}\}_{k=1}^{K})$
9:  Broadcast the global model $w_{t+1}$ to all clients
10:end for
11:ClientUpdate($\mathcal{D}_{k},w$):
12:Split $\mathcal{D}_{k}$ into batches $\{b\}$
13:for $e=0$ to $E$ do $\triangleright$ $e$-th local epoch
14:  for each batch $b$ in $\{b\}$ do
15:   Extract inputs $(x,y)$ from $b$
16:   $\mathbf{f}\leftarrow\text{FeatureExtractor}(w;x)$
17:   $\mathbf{H}\leftarrow\text{LADM}(\mathbf{M},\mathbf{f})$ $\triangleright$ Eq. 3
18:   Compute loss $\mathcal{L}_{\text{total}}(w;\mathbf{M},\mathbf{H})$ $\triangleright$ Eqs. 5789
19:   $w\leftarrow w-\eta\nabla_{w}\mathcal{L}_{\text{total}}$
20:  end for
21:end for
22:Return $w$

### v2

@@ Introduction
We consider a multi-label, label-skewed FL setting with $K$ clients collaboratively training a shared global model. Each client $k$ holds a private dataset $\mathcal{D}_{k}=\{(x_{i}^{k},y_{i}^{k})\}_{i=1}^{N_{k}}$ of size $N_{k}$, where the input image $x_{i}^{k}\in\mathbb{R}^{H\times W\times 3}$ and the label $y_{i}^{k}\in\{0,1\}^{C}$ is a multi-hot vector over the $C$ classes. Let $\mathcal{C}=\{1,\dots,C\}$ denote the global class index set. Due to label skew, client $k$ may only observe a subset $\mathcal{C}_{k}\subseteq\mathcal{C}$, and even for shared classes, the class-conditional distributions can differ across clients. The goal is to learn a single global model that accurately recognizes all classes in $\mathcal{C}$ without sharing raw data among clients.
Neural Collapse (NC)
@@
Simplex Equiangular Tight Frame (ETF).
A simplex ETF matrix $\mathbf{M}=[\mathbf{m}_{c}]_{c=1}^{C}\in\mathbb{R}^{d\times C}$ composed of $C$ column vectors, each corresponding to a class prototype in $\mathbf{m}_{c}\in\mathbb{R}^{d}$. A standard construction is:
$\mathbf{M}\;=\;\sqrt{\frac{C}{C-1}}\;\mathbf{U}\;\Big(\mathbf{I}_{C}-\frac{1}{C}\,\mathbf{1}_{C}\mathbf{1}_{C}^{\top}\Big),$ (1)
where $\mathbf{U}\in\mathbb{R}^{d\times C}$ denotes a rotation orthogonal matrix $(\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}_{C})$. $\mathbf{I}_{C}$ is the $C\times C$ identity matrix. $\mathbf{1}_{C}$ is the $C$-dimensional all-ones vector.
This yields unit-norm columns with equal pairwise inner products:
$\mathbf{m}_{a}^{\top}\mathbf{m}_{b}=\begin{cases}1,&a=b,\\[2.0pt] -\dfrac{1}{C-1},&a\neq b,\end{cases}\qquad a,b\in\mathcal{C}.$ (2)
Hence, all prototypes have equal $\ell_{2}$ norm and identical pairwise angles, forming a centered regular simplex.
Neural Collapse properties.
At the end of training, over-parameterized networks exhibit the NC phenomenon. Empirically, the following regularities are observed:
  • Variability Collapse ($\mathcal{NC}_{1}$). Within-class feature variance collapses: features of the same class concentrate at their class mean.
  • Convergence to Simplex ETF ($\mathcal{NC}_{2}$). Class means arrange as the vertices of a simplex ETF, forming a maximally symmetric and equidistant configuration.
  • Self-Duality ($\mathcal{NC}_{3}$). Upon appropriate rescaling, the final-layer classifier weights align with the class means, sharing the same simplex ETF geometry.
@@
The NC phenomenon suggests a promising structure for optimal feature clustering in the latent feature space under well-trained conditions. We leverage this geometry to shape the latent feature space in the challenging multi-label FL setting, where client data often deviates from the global distribution. An overview of our framework is shown in Figure 2.
Label-Aware Disentanglement Module (LADM)
Multi-label images may contain multiple object categories, making a single pooled representation suboptimal because it compresses multiple semantics into one vector. Inspired by recent advances in multi-label learning (Liu et al.,, 2021; Ridnik et al.,, 2023; Xu et al.,, 2023), we incorporate an attention mechanism to extract class-specific information from the image-level representation, thereby enhancing the discriminative power of each label.
[…]
$\mathbf{h}_{ic}\;=\;\operatorname{MultiHeadAttn}(\mathbf{m}_{c},\;\mathbf{f}_{i}),\quad c\in\mathcal{C}$ (3)
and stacks them into
$\mathbf{H}_{i}\;=\;\begin{bmatrix}\mathbf{h}_{i1}^{\top}\\[-2.0pt] \vdots\\[-2.0pt] \mathbf{h}_{iC}^{\top}\end{bmatrix}\in\mathbb{R}^{C\times d}.$ (4)
Sharing the fixed query matrix $\mathbf{M}$ across clients mitigates aggregation conflicts and encourages globally aligned, disentangled features.
@@
After disentangling image-level features into class-wise representations with LADM, each sample $i\in[N_{k}]$ yields a feature $\mathbf{h}_{ic}\in\mathbb{R}^{d}$ for every class $c\in\mathcal{C}$. To avoid client-specific classifier drift caused by label imbalance or missing classes, we fix the classifier to the global simplex ETF used as the LADM query matrix, $\mathbf{M}=[\mathbf{m}_{c}]_{c=1}^{C}\in\mathbb{R}^{d\times C}$. Given $\mathbf{h}_{ic}$ and its corresponding prototype $\mathbf{m}_{c}$, we compute a class logit and pass it through the sigmoid to obtain a binary prediction:
$\hat{y}_{ic}\;=\;\sigma\!\big(\mathbf{h}_{ic}^{\top}\mathbf{m}_{c}\big),$ (5)
and train with a binary cross-entropy (BCE) objective
$\mathcal{L}_{\text{BCE}}=-\frac{1}{N}\frac{1}{C}\sum_{i=1}^{N}\sum_{c=1}^{C}\Big[y_{ic}\,\log\hat{y}_{ic}+(1-y_{ic})\,\log(1-\hat{y}_{ic})\Big].$ (6)
Here, $y_{ic}\in\{0,1\}$ is the ground-truth multi-label indicator, $\sigma(\cdot)$ denotes the sigmoid. This fixed classifier improves class separability and reduces divergence between local and global label distributions.
@@
Each sample produces $C$ class-wise features $\{\mathbf{h}_{ic}\}_{c=1}^{C}$. Let the positive and negative class sets for sample $i$ be $\mathcal{C}_{i}^{+}=\{\,c\mid y_{ic}=1\,\}$ and $\mathcal{C}_{i}^{-}=\{\,c\mid y_{ic}=0\,\}$, respectively. We introduce two regularizers: one to suppress noise from negative features and another to encourage compact clustering of positive features.
Negative Feature Rejection Loss.
While $\mathcal{L}_{\text{BCE}}$ discourages alignment of a negative feature $\mathbf{h}_{ic}$ with its own prototype $\mathbf{m}_{c}$ when $y_{ic}=0$, it does not prevent $\mathbf{h}_{ic}$ from spuriously aligning with other prototypes. We therefore penalize high similarity between every negative feature and all non-self prototypes:
$\hat{s}_{icr}\;=\;\sigma\!\big(\mathbf{h}_{ic}^{\top}\mathbf{m}_{r}\big),\quad c\in\mathcal{C}_{i}^{-},\;r\in\mathcal{C}\setminus\{c\},$ (7)
$\mathcal{L}_{\text{Neg}}=-\,\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|\mathcal{C}_{i}^{-}|}\sum_{c\in\mathcal{C}_{i}^{-}}\frac{1}{C-1}\sum_{\begin{subarray}{c}r=1\\ r\neq c\end{subarray}}^{C}\mathbb{I}\!\big(\hat{s}_{icr}>\tau\big)\,\log\!\big(1-\hat{s}_{icr}\big).$ (8)
The indicator $\mathbb{I}(\cdot)$ filters out low-similarity pairs so that only confident negatives contribute. In our experiments, we set $\tau=0.3$.
Positive Feature Contrastive Loss.
@@
To encourage compact and discriminative class-wise clustering in the latent feature space, we introduce a contrastive loss. This loss enforces each positive feature $\mathbf{h}_{ic}$ to be closer to its own prototype than to others via a prototype-softmax:
$\mathcal{L}_{\text{Pos}}=-\,\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|\mathcal{C}_{i}^{+}|}\sum_{c\in\mathcal{C}_{i}^{+}}\log\frac{\exp\!\big(\mathbf{h}_{ic}^{\top}\mathbf{m}_{c}\big)}{\sum_{r=1}^{C}\exp\!\big(\mathbf{h}_{ic}^{\top}\mathbf{m}_{r}\big)}.$ (9)
Total Objective.
The total loss can be expressed as:
$\mathcal{L}_{\text{total}}\;=\;\mathcal{L}_{\text{BCE}}+\lambda_{1}\,\mathcal{L}_{\text{Neg}}+\lambda_{2}\,\mathcal{L}_{\text{Pos}}.$ (10)
where $\lambda_{1},\lambda_{2}\geq 0$ balance the contributions of the regularizers.
FedNCAlign-ML
We propose FedNCAlign-ML, a framework that enhances the standard FL protocol with an NC-inspired local training strategy tailored for multi-label tasks and skewed local data distributions. In this framework, each client first encodes its input images into features, which are then disentangled into class-specific embeddings using LADM. These embeddings are aligned with a shared simplex ETF matrix that serves simultaneously as fixed classifier prototypes and as class-wise queries for LADM, thereby enforcing a consistent global geometry across clients despite label imbalance. To further improve representation quality, we incorporate two regularizers: (i) a negative rejection loss to suppress spurious alignments of irrelevant features, and (ii) a positive contrastive loss to encourage compact intra-class clustering and clear inter-class separation. The novelty of FedNCAlign-ML lies in this local training design, which integrates disentangled class-wise embeddings, ETF-based alignment, and geometry-aware regularization. The complete procedure is summarized in Algorithm 1.
Algorithm 1 FedNCAlign-ML: Federated Neural Collapse Alignment for Multi-Label Learning
Input: $K$ clients with datasets $\{\mathcal{D}_{k}\}_{k=1}^{K}$; initial global model $w_{0}$; predefined ETF matrix $\mathbf{M}$; learning rate $\eta$; local epochs $E$; communication rounds $T$.
1:Server executes:
2:Initialize $w\leftarrow w_{0}$
3:for $t=0$ to T1T-1 do $\triangleright$ communication rounds
4:  for each client $k\in\{1,\dots,K\}$ in parallel do
5:   $w_{t+1}^{k}\leftarrow\textsc{ClientUpdate}(\mathcal{D}_{k},w_{t},\mathbf{M})$
6:  end for
7:  $w_{t+1}\leftarrow\textsc{ModelAggregation}(\{w_{t+1}^{k}\}_{k=1}^{K})$
8:  Broadcast $w_{t+1}$ to all clients
9:end for
10:
11:function ClientUpdate($\mathcal{D}_{k},w,\mathbf{M}$)
12:  for $e=1$ to $E$ do $\triangleright$ local epochs
13:   for each batch $(x,y)\subset\mathcal{D}_{k}$ do
14:     $\mathbf{f}\leftarrow\textsc{FeatureExtractor}(w,x)$
15:     $\mathbf{H}\leftarrow\textsc{LADM}(\mathbf{f},\mathbf{M})$
16:     Compute prediction $\hat{y}_{c}=\sigma(\mathbf{h}_{c}^{\top}\mathbf{m}_{c})$ $\triangleright$ Eqs. 5
[…]
18:     Update $w\leftarrow w-\eta\nabla_{w}\mathcal{L}_{\text{total}}$
19:   end for
20:  end for
21:  return $w$
22:end function
@@
(a) FC classifier
(b) LADM + fixed ETF classifier
(c) LADM + ETF + Dual regularization